OpenStudy (blurryface):

The names of 9 boys and 2 girls are in a hat. What is the probability of first drawing a girl’s name and then a boy’s name from the hat?

OpenStudy (blurryface):

hello @Kinged

OpenStudy (kinged):

What's the probability of fist drawing a girl's name?

OpenStudy (blurryface):

idk

OpenStudy (kinged):

It's just the number of girls divided by the total number of names in the hat.

OpenStudy (blurryface):

that is not correct

OpenStudy (kinged):

That's just the first step of the answer.

OpenStudy (blurryface):

that would be 5.5 or 11/2 and that is not correct

OpenStudy (blurryface):

oh silly me

OpenStudy (kinged):

First of all, You mean 2/11 not 11/2.

OpenStudy (kinged):

Now, you have to get the probability of picking a boy's name thereafter.

OpenStudy (blurryface):

2/11 girls and 9/11 boys

OpenStudy (blurryface):

what next?

OpenStudy (kinged):

No. For the boys, the total has changed. By how much?

OpenStudy (blurryface):

7/11?

OpenStudy (kinged):

They are still 9 boys, only we have removed a person from the total. After we calculated the probability of pick a girl's name.

OpenStudy (blurryface):

What is the equation we will use?

OpenStudy (kinged):

P(girl then boy)=P(girl)*P(boy)

OpenStudy (blurryface):

I don't understand

OpenStudy (sshayer):

when you draw name of one girl ,you are left with 10 names

OpenStudy (sshayer):

9 boys name and one girl name.

OpenStudy (blurryface):

now what?

OpenStudy (kinged):

The probability of picking a boy's name is the total number of boys divided by the total number of name minus one.

OpenStudy (kinged):

When you do this, return to the equation I stated previously.

OpenStudy (blurryface):

Could you do it for me?

OpenStudy (kinged):

Nope. Sorry, though.

OpenStudy (blurryface):

do you not know how to do it?

OpenStudy (kinged):

I know how to do it, just not for you.

OpenStudy (kinged):

I outlined the steps as clear as I could. Just revise them.

OpenStudy (blurryface):

boys = 8/11? girls 2/11? what next

OpenStudy (kinged):

I'm sorry, but that's the code of ethics of this website. You have to understand.

OpenStudy (blurryface):

I learn by example...

OpenStudy (kinged):

The boys are still nine. The total is what's changed.

OpenStudy (kinged):

Since you just picked a person before.

OpenStudy (blurryface):

the answer is 9/55... I understand it now

OpenStudy (blurryface):

is that correct?

OpenStudy (sshayer):

\[\frac{ 2 }{ 11 }\times \frac{ 9 }{ 10 }=?\]

OpenStudy (kinged):

YEPPP.

OpenStudy (blurryface):

I got 1 more question